OpenStudy (anonymous):

Simplify using the rules of exponents (3x^-3y^-5)(-4x^5y^-2)

OpenStudy (rulnick):

-12 x^2 y^-7

OpenStudy (anonymous):

Ohh how did you get that answer and dont you need to put the -7 exponent on the bottom?

OpenStudy (rulnick):

Since it's all multiplication, you can group the like variables: 3(-4) = -12 x^-3 x^5 = x^2 y^-5 y^-2 = y^-7

OpenStudy (rulnick):

And, yes, -12 x^2 y^-7 can (but need not) be written -12 x^2 / y^7.

OpenStudy (anonymous):

OOh ok thank you :)

OpenStudy (anonymous):

Oh wait what about this one.... 36x^-7y^5z^-2/4x^4y^-3z^-4 Would you combine the like terms first or put the negative exponents on the bottom to the top first and then combine like terms